• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C.E.Hyde-Wright

    Log entry time 23:54:05 on January24,2002

    Entry number 80642

    keyword=Shift Summary

    24-January-2002
    Shift Summary
    16:45 Beam restore 5 micro Amp
    check beam spot on BeO
    move beam to BPM-B-x=-2.0 mm
    BPM-B-y=+1.5 mm
    harp scan run 1482
    turn on raster: request 5mm to obtain 2mm (+/- 1mm)
    17:15 move to cryo target
    cryo target motion frozen.
    18:30 10 micro A on LH2
    efficiency runs 1605... with only S0: ps7=1
    19:00 lower beam to 5 micro A
    19:15 Start Kinematics 2A
    move to optics target
    19:35 start optics target run 1610
    19:50 MCC tests 20min
    20:10 15 cm LH2 magnet-off on-endpoint run 1611
    20:30 15 cm LH2 magnet-500A on-endpoint run 1612
    21:00 moving target to radiator in. Target motion frozen
    very difficult to get Bogie motor on to move spec-L
    angle 35.678 deg mark 35.5 vernier 62.0
    21:45 15 cm LH2+rad magnet-500A off-endpoint run 1613
    raise calorimeter trigger threshold
    23:00 15 cm LH2+rad magnet-off off-endpoint run 1620
    23:35 15cm LH2 magnet-500A off-endpoint run 1623
    *****************
    4.75 Hrs Production Running !!!!!!!!!!!!!!!!!!
    *****************